# ConRFT: A Reinforced Fine-tuning Method for VLA Models via Consistency Policy

### v1

@@ I Introduction
Recent advancements in training generalist robotic policies using Vision-Language-Action (VLA) models have demonstrated remarkable capabilities in understanding and executing various manipulation tasks. These successes are primarily attributed to large-scale imitation-style pre-training and grounding with robot actions [1, 2, 3, 4, 5, 6]. While pre-trained policies capture powerful representations, they often fall short when handling the complexities of real-world scenarios [7]. Fine-tuning with domain-specific data is essential to optimize model performance for downstream tasks [8, 6]. While Supervised Fine-Tuning (SFT) of the VLA model using human teleoperation data remains the predominant adaptation approach, this process faces significant challenges: the model’s performance heavily relies on the quality and quantity of task-specific data. However, these human-collected teleoperation datasets may not consistently provide optimal trajectories due to inherent issues such as sub-optimal demonstration and inconsistent action [9].
Significant progress in Large-Language-Models (LLMs) and Vision-Language-Models (VLMs) have highlighted the value of reinforcement learning as a powerful tool for bridging the gap between policy capabilities and human preference [10, 11, 12, 13] or improving model reasoning [14]. In addition, deploying reinforcement learning (RL) with task-specific reward functions to learn from online interaction data is also a promising direction [15, 16, 17]. However, extending these insights to VLA models presents unique challenges because, unlike LLMs, VLA models necessitate direct physical interaction in real-world robotic manipulation tasks. The safety and cost constraints of collecting data in contact-rich environments demand high sample efficiency and risk-aware exploration, making a straightforward implementation of RL infeasible. Recent work has attempted to leverage RL to address the challenges faced in SFT [9, 18]; these methods primarily focus on utilizing RL for data augmentation or quality improvement rather than directly optimizing VLA models through RL objectives. This reliance limits the policy’s ability to explore states out of the demonstration dataset, thus undermining the potential benefits of RL-based fine-tuning in real-world settings.
@@
    We present a Consistency-based Reinforced Fine-Tuning (ConRFT) method, a unified pipeline with the same training objective both for offline and online fine-tuning.
@@ II-A Reinforced Fine-tuning for Large Models
RL has been widely adopted for fine-tuning LLMs and VLMs. Early works have primarily focused on RL incorporating human feedback [10, 25, 11, 12, 13] by learning from human preferences or by integrating task-specific rewards without explicit human preference [15, 16, 17, 26]. While many of these approaches employ on-policy algorithms (e.g., PPO [27]) to fine-tune pre-trained policies [28, 16, 29], they typically demand large amounts of interaction data to achieve desirable performance [30]. While RL has demonstrated success in many domains, it typically learns within self-generated synthetic environments rather than real-world environments. This gap prevents direct transfer for VLA models, which require real-world interaction. Our work addresses this discrepancy by developing RL frameworks tailored for efficient real-world VLA fine-tuning.
@@ II-C Offline-to-online Methods
Offline-to-online RL aims to leverage offline datasets to initialize a policy, which is then fine-tuned via online interactions for improved sample efficiency [42, 19]. Existing works commonly adopt an offline pre-training stage followed by an online fine-tuning stage [43, 19, 44, 20], mixing offline and online data as training proceeds. This offline-to-online pipeline is similar to our proposed two-stage fine-tuning approach that exploits pre-collected data to bootstrap policy training and then fine-tunes the policy in the real-world tasks [34]. Most offline-to-online methods assume the availability of large-scale, diverse datasets with sufficient state coverage [45, 46], a condition rarely met in real-world deployments. We explore leveraging pre-trained VLA models as the base policy to enable sample-efficient policy refinement, achieving superior fine-tuning performance even under stringent demonstration data constraints.
@@ III Problem Setup and Preliminaries
Since demonstrations are often limited, inconsistent, and sub-optimal, preventing the policy from covering diverse states, SFT struggles in real-world, contact-rich robotic tasks. To address these issues, we formulate each robotic task as a Markov Decision Process (MDP), where the goal of RL is to find the optimal policy in the MDP, $\mathcal{M}=(\mathcal{S},\mathcal{A},\mathcal{P},r,\rho,\gamma)$, where $s\in\mathcal{S}$ denotes the state space and $a\in\mathcal{A}$ denotes the action space. $\mathcal{P}(s^{\prime}|s,a)$ is the environmental transition probabilities that depend on the system dynamics, and $\rho(s)$ denotes the initial state distribution. $r(s,a)$ and $\gamma\in(0,1)$ are the reward functions and the reward discount factor. The policy $\pi$ is estimated by maximizing the cumulative expected value of the reward, denoted as $V^{\pi}(s)=\mathbb{E}_{\pi}[\sum_{t=0}^{H}\gamma^{t}r(s_{t},a_{t})|s_{0}=s,a_{t}\sim\pi(s_{t}),s_{t+1}\sim p(\cdot|s_{t},a_{t})]$. The Q-function of a given policy $\pi$ is denoted as $Q^{\pi}(s,a)=\mathbb{E}_{\pi}[\sum_{t=0}^{H}\gamma^{t}r(s_{t},a_{t})|s_{0}=s,a_{0}=a,a_{t+1}\sim\pi(s_{t+1}),s_{t+1}\sim p(\cdot|s_{t},a_{t})]$. By coupling the VLA policy with the learned Q-function, RFT allows the VLA model to refine its behavior based on trial-and-error interactions and task-specific feedback.
Figure 1: Overview of ConRFT. This figure illustrates the architecture of our reinforced fine-tuning approach, which comprises two stages: the offline stage (Cal-ConRFT) and the online stage (HIL-ConRFT). Our method is built on top of a pre-trained VLA model. Both stages use a unified consistency-based training objective to fine-tune the VLA model. During the offline stage, we only use pre-collected demonstrations for training. During the online stage, a human operator can intervene in the robot policy via teleoperation tools, such as a SpaceMouse. We use both pre-collected demonstrations, policy transitions, and human interventions for online fine-tuning via HIL-ConRFT.
@@ IV Method
The proposed unified pipline ConRFT consists of two stages: offline fine-tuning followed by online fine-tuning to optimize robotic policies. The overview of our method is shown in Fig. 1. In the following sections, we provide a detailed description of the two stages. And we provide the whole pipeline illustration in Appendix -A.
@@ IV-A Stage I: Offline Fine-tuning with Cal-ConRFT
where $Q_{\theta}$ is the learned Q-function parameterized by $\theta$, $\overline{Q}_{\overline{\theta}}$ is the delayed target Q-function parameterized by $\overline{\theta}$. $\mathcal{B}^{\pi}\overline{Q}(s,a)=r(s,a)+\gamma\mathbb{E}_{a^{\prime}\sim\pi(\cdot|s^{\prime})}(\overline{Q}(s^{\prime},a^{\prime}))$ is the backup operator. $\alpha$ is a hyper-parameter to control the conservative penalty. And $\mathcal{D}$ is the demo buffer that stores demonstrations.
@@
To address this issue, we propose augmenting the offline training process by incorporating a BC loss. The BC loss directly minimizes the difference between the actions generated by the policy and those from the demonstrations. By incorporating BC loss, we encourage the model to imitate the behaviors from the demonstrations, providing additional supervisory signals during the offline stage. This helps the VLA model to learn a more effective policy and initialize a stable Q function with few demonstrations, especially in the case of contact-rich manipulation tasks where precision is critical.
Motivated by combining the BC loss with Q guidance under a consistency-based objective [22], we introduce Cal-ConRFT in the offline stage. This approach employs a consistency policy as the action head for fine-tuning the VLA model, addressing two key concerns: 1) it helps leverage inconsistent and sub-optimal demonstrations that often arise in pre-collected data, and 2) compared to diffusion-based action head, the consistency-based action head remains computationally lightweight for efficient inference [22, 47]. The consistency policy is a diffusion-model-based policy that learns to map random actions sampled from the unit Gaussian to generate actions drawn from the expert action distribution conditioned on the current state. For the consistency policy, we discretize the diffusion horizon $[\epsilon,K]$ into $M$ sub-intervals with boundaries $k_{1}=\epsilon\leq k_{2}\leq\cdots\leq k_{m}=K$ and $\epsilon=0.002$. Specifically, the VLA model with a consistency policy as the action head is given by:
@@
$\mathcal{L}_{\pi}^{offline}(\psi)=-\eta\mathbb{E}_{s\sim\mathcal{D},a\sim\pi_{\psi}}[Q(s,a)]$ (3)
$+$ $\beta\mathbb{E}_{(s,a)\sim\mathcal{D},m\sim\mathcal{U}[1,M-1]}[d(f_{\psi}(a+k_{m}z,k_{m}|E(s)),a)]$
where $z\sim\mathcal{N}(0,I)$, $d$ stands for the Euclidean distance $d(x,y)=||x-y||_{2}$, $\beta$ and $\eta$ are two hyper-parameters to balance the BC loss and Q loss. This combination enables efficient policy learning and stable value estimation, even with a small set of demonstrations, by aligning value estimates with expert actions and improving policy performance during offline training. Moreover, it provides a reliable initialization for the online stage, facilitating safe and effective exploration.
@@
Pick Banana 45 40 50 60 (+50%) 80% (+100%) 90 (+80%) 63.7 57.8 67.5 (0.9x) 56.1 (1.1x) 51.2 (1.1x)
Put Spoon 45 50 55 90 (+80%) 80% (+60%) 100 (+82%) 49.9 57.2 50.9 (1.0x) 45.3 (1.1x) 22.6 (2.5x)
Open Drawer 15 35 30 80 (+129%) 60% (+71%) 100 (+233%) 63.6 61.7 48.4 (1.3x) 57.1 (1.1x) 32.4 (1.8x)
Pick Bread 45 65 55 65 (+0%) 80% (+23%) 100 (+82%) 53.2 49.1 65.6 (0.8x) 51.7 (1.0x) 31.6 (1.6x)
Open Toaster 30 30 30 75 (+116%) 100% (+233%) 100 (+233%) 51.2 50.7 43.4 (1.2x) 34.3 (1.5x) 22.1 (2.3x)
Put Bread 60 5 20 60 (+1100%) 75% (+1400%) 100 (+400%) 102 84.8 74.2 (1.4x) 72.1 (1.4x) 36.6 (2.3x)
Insert Wheel 60 35 35 40 (+14%) 30% (-14%) 80 (+129%) 42.7 43.4 53.0 (0.8x) 47.4 (0.9x) 21.9 (2.0x)
Hang Chinese Knot 90 55 40 50 (-10%) 65% (+18%) 100 (+150%) 52.6 54.9 47.5 (1.1x) 44.4 (1.3x) 26.8 (2.0x)
Average 48.8 39.4 39.4 65 (+65%) 71.3% (+81%) 96.3 (+144%) 59.9 57.5 56.3 (1.1x) 51.1 (1.2x) 30.7 (1.9x)
@@ IV-B Stage II: Online Fine-tuning with HIL-ConRFT
$\mathcal{L}_{\pi}^{online}(\psi)=-\eta\mathbb{E}_{s\sim(\mathcal{D}\cup\mathcal{R}),a\sim\pi_{\psi}}[Q(s,a)]$ (5)
$+$ $\beta\mathbb{E}_{(s,a)\sim(\mathcal{D}\cup\mathcal{R}),m\sim\mathcal{U}[1,M-1]}[d(f_{\psi}(a+k_{m}z,k_{m}|E(s)),a)]$
Note that this objective closely mirrors Equation 3 from the offline stage, enabling a quick adaption to online fine-tuning.
@@ VI Limitations
Although our approach demonstrates strong performance and sample efficiency for fine-tuning VLA models in real-world manipulation tasks, several limitations remain:
@@ VI-A Sensitivity to Reward Engineering
In this work, we implement a task-specific binary classifier to calculate the reward for RL. However, the inherent distributional shift between the classifier’s training data and the state-action distributions generated during RL exploration creates a critical vulnerability ,as it can lead the learned policy to engage in reward hacking, exploiting unintended behaviors where the classifier provides inaccurate rewards. For instance, the robot might position its end-effector at a specific location that triggers a false positive, causing the policy to converge to an incorrect behavior. Moreover, because these reward classifiers typically provide only sparse feedback, the policy may learn slowly, even with the help of human interventions. Introducing dense reward signals could enhance sample efficiency and accelerate policy convergence, but it would also require more sophisticated reward engineering for real-world environments.

### v2

@@ I Introduction
Recent advancements in training generalist robotic policies using Vision-Language-Action (VLA) models have demonstrated remarkable capabilities in understanding and executing various manipulation tasks. These successes are primarily attributed to large-scale imitation-style pre-training and grounding with robot actions [1, 2, 3]. While pre-trained policies capture powerful representations, they often fall short when handling the complexities of real-world scenarios [4]. Fine-tuning with domain-specific data is essential to optimize model performance for downstream tasks [3, 5]. While Supervised Fine-Tuning (SFT) of the VLA model using human teleoperation data remains the predominant adaptation approach, this process faces significant challenges: the model’s performance heavily relies on the quality and quantity of task-specific data. However, these human-collected datasets may not consistently provide optimal trajectories due to inherent issues such as sub-optimal data and inconsistent action [6].
Significant progress in Large-Language-Models (LLMs) and Vision-Language-Models (VLMs) have highlighted the value of reinforcement learning as a powerful tool for bridging the gap between policy capabilities and human preference [7, 8, 9] or improving model reasoning [10]. In addition, deploying reinforcement learning (RL) with task-specific reward functions to learn from online interaction data is also a promising direction [11, 12, 13]. However, extending these insights to VLA models presents unique challenges because, unlike LLMs, VLA models necessitate direct physical interaction in real-world robotic tasks. The safety and cost constraints of collecting data in contact-rich environments demand high sample efficiency and risk-aware exploration, making a straightforward implementation of RL infeasible. Recent work has attempted to leverage RL to address the challenges faced in SFT [6, 14], while these methods primarily focus on utilizing RL for data augmentation or quality improvement rather than directly optimizing VLA models through RL objectives. This limits the policy’s ability to explore states out of the demonstration dataset, thus undermining the potential benefits of RL-based fine-tuning in real-world settings.
@@
    We present a Consistency-based Reinforced Fine-Tuning (ConRFT) approach, a novel pipeline with the unified training objective both for offline and online fine-tuning.
@@ II-A Reinforced Fine-tuning for Large Models
RL has been widely adopted for fine-tuning LLMs and VLMs. Early works have primarily focused on RL incorporating human feedback [7, 8, 9, 21, 22] by learning from human preferences or by integrating task-specific rewards without explicit human preference [11, 12, 13, 23]. While many of these approaches employ on-policy algorithms (e.g., PPO [24]) to fine-tune pre-trained policies [12, 25, 26], they typically demand large amounts of interaction data to achieve desirable performance [27, 28]. While RL has demonstrated success in many domains, it typically learns within self-generated synthetic environments rather than real-world environments. This gap prevents direct transfer for VLA models, which require real-world interaction. Our work addresses this discrepancy by developing RL frameworks tailored for efficient real-world VLA fine-tuning.
@@ II-C Offline-to-online Methods
Offline-to-online RL aims to leverage offline datasets to initialize a policy, which is then fine-tuned via online interactions for improved sample efficiency [15]. Existing works commonly adopt an offline pre-training stage followed by an online fine-tuning stage [15, 40, 41, 16], mixing offline and online data as training proceeds. This offline-to-online pipeline is similar to our proposed two-stage fine-tuning approach that exploits pre-collected data to bootstrap policy training and then fine-tunes the policy in the real-world tasks [32]. Most offline-to-online methods assume the availability of large-scale, diverse datasets with sufficient state coverage [42, 43], a condition rarely met in real-world deployments. We explore leveraging pre-trained VLA models as the base policy to enable sample-efficient policy refinement, achieving superior fine-tuning performance even under stringent demonstration data constraints.
@@ III Problem Setup and Preliminaries
Since demonstrations are often limited, inconsistent, and sub-optimal, preventing the policy from covering diverse states, SFT struggles in real-world, contact-rich robotic tasks. To address these issues, we formulate each robotic task as a Markov Decision Process (MDP), where the goal of RL is to find the optimal policy in the MDP, $\mathcal{M}=(\mathcal{S},\mathcal{A},\mathcal{P},r,\rho,\gamma)$, where $s\in\mathcal{S}$ denotes the state space and $a\in\mathcal{A}$ denotes the action space. $\mathcal{P}(s^{\prime}|s,a)$ is the environmental transition probabilities that depend on the system dynamics, and $\rho(s)$ denotes the initial state distribution. $r(s,a)$ and $\gamma\in(0,1)$ are the reward function and the reward discount factor. The policy $\pi$ is estimated by maximizing the cumulative expected value of the reward, denoted as $V^{\pi}(s)=\mathbb{E}_{\pi}[\sum_{t=0}^{H}\gamma^{t}r(s_{t},a_{t})|s_{0}=s,a_{t}\sim\pi(s_{t}),s_{t+1}\sim p(\cdot|s_{t},a_{t})]$. The Q-function of a given policy $\pi$ is denoted as $Q^{\pi}(s,a)=\mathbb{E}_{\pi}[\sum_{t=0}^{H}\gamma^{t}r(s_{t},a_{t})|s_{0}=s,a_{0}=a,s_{t+1}\sim p(\cdot|s_{t},a_{t})]$. $H$ represents the maximum episode step of a trajectory. By coupling the VLA policy with the learned Q-function, RFT allows the VLA model to refine its behavior based on trial-and-error interactions and task-specific feedback.
Figure 1: Overview of ConRFT. This figure illustrates the architecture of our reinforced fine-tuning approach for a pre-trained VLA model, which comprises two stages: the offline Cal-ConRFT and the online HIL-ConRFT. Both stages use a unified consistency-based training objective. During the offline stage, we only use pre-collected demonstrations for fine-tuning. During the online stage, a human operator can intervene in the robot policy via teleoperation tools(e.g. a SpaceMouse). And we use both pre-collected demonstrations, policy transitions, and human interventions for fine-tuning.
@@ IV Method
The proposed pipline ConRFT consists of two stages: offline fine-tuning followed by online fine-tuning to optimize robotic policies, as shown in Fig. 1. In the following sections, we provide a detailed description of the two stages, with the pipeline illustration in Appendix -A.
@@ IV-A Stage I: Offline Fine-tuning with Cal-ConRFT
where $Q_{\theta}$ is the learned Q-function parameterized by $\theta$, $\overline{Q}_{\overline{\theta}}$ is the delayed target Q-function parameterized by $\overline{\theta}$. $\mathcal{B}^{\pi}\overline{Q}(s,a)=r(s,a)+\gamma\mathbb{E}_{a^{\prime}\sim\pi(\cdot|s^{\prime})}(\overline{Q}(s^{\prime},a^{\prime}))$ is the Bellman backup operator. $\alpha$ is a hyper-parameter to control the conservative penalty. And $\mathcal{D}$ is the demo buffer that stores demonstrations.
@@
To address this issue, we propose augmenting the offline training process by incorporating a BC loss. The BC loss directly minimizes the difference between the actions generated by the policy and those from the demonstrations. By incorporating BC loss, we encourage the model to imitate the behaviors from the demonstrations, providing additional supervisory signals during the offline stage. This helps the VLA model to learn a more effective policy and initialize a stable Q function with few demonstrations, especially in the case of contact-rich manipulation tasks where control precision is critical.
Motivated by combining the BC loss with Q guidance under a consistency-based objective [18], we introduce Cal-ConRFT in the offline stage. This approach employs a consistency policy as the action head for fine-tuning the VLA model, addressing two key concerns: 1) it helps leverage inconsistent and sub-optimal demonstrations that often arise in pre-collected data, and 2) compared to diffusion-based action head, the consistency-based action head remains computationally lightweight for efficient inference [18, 44, 45]. The consistency policy is a diffusion-model-based policy [46] that learns to map random actions sampled from the unit Gaussian to generate actions drawn from the expert action distribution conditioned on the current state. For the consistency policy, we discretize the diffusion horizon $[\epsilon,K]$ into $M$ sub-intervals with boundaries $k_{1}=\epsilon\leq k_{2}\leq\cdots\leq k_{m}=K$ and $\epsilon=0.002$. Specifically, the VLA model with a consistency policy as the action head is given by:
@@
$\mathcal{L}_{\pi}^{offline}(\psi)=\beta\mathcal{L}_{\pi}^{BC}+\eta\mathcal{L}_{\pi}^{Q}$ (3)
where BC loss $\mathcal{L}_{\pi}^{BC}=\mathbb{E}_{(s,a)\sim\mathcal{D},m\sim\mathcal{U}[1,M-1]}[d(f_{\psi}(a+k_{m}z,k_{m}|E(s)),a)]$, $z\sim\mathcal{N}(0,I)$, $d$ stands for the Euclidean distance $d(x,y)=||x-y||_{2}$, and Q loss $\mathcal{L}_{\pi}^{Q}=-\mathbb{E}_{s\sim\mathcal{D},a\sim\pi_{\psi}}[Q(s,a)]$. $\beta$ and $\eta$ are two hyper-parameters to balance the BC loss and Q loss. This combination enables efficient policy learning and stable value estimation, even with a small set of demonstrations, by aligning value estimates with expert actions and improving policy performance during offline training. Moreover, it provides a reliable initialization for the online stage, facilitating safe and effective exploration.
@@
Pick Banana 45 40 50 60 (+50%) 80 (+100%) 90 (+80%) 63.7 57.8 67.5 (0.9x) 56.1 (1.1x) 51.2 (1.1x)
Put Spoon 45 50 55 90 (+80%) 80 (+60%) 100 (+82%) 49.9 57.2 50.9 (1.0x) 45.3 (1.1x) 22.6 (2.5x)
Open Drawer 15 35 30 80 (+129%) 60 (+71%) 100 (+233%) 63.6 61.7 48.4 (1.3x) 57.1 (1.1x) 32.4 (1.8x)
Pick Bread 45 65 55 65 (+0%) 80 (+23%) 100 (+82%) 53.2 49.1 65.6 (0.8x) 51.7 (1.0x) 31.6 (1.6x)
Open Toaster 30 30 30 75 (+116%) 100 (+233%) 100 (+233%) 51.2 50.7 43.4 (1.2x) 34.3 (1.5x) 22.1 (2.3x)
Put Bread 60 5 20 60 (+1100%) 75 (+1400%) 100 (+400%) 102 84.8 74.2 (1.4x) 72.1 (1.4x) 36.6 (2.3x)
Insert Wheel 60 35 35 40 (+14%) 30 (-14%) 80 (+129%) 42.7 43.4 53.0 (0.8x) 47.4 (0.9x) 21.9 (2.0x)
Hang Chinese Knot 90 55 40 50 (-10%) 65 (+18%) 100 (+150%) 52.6 54.9 47.5 (1.1x) 44.4 (1.3x) 26.8 (2.0x)
Average 48.8 39.4 39.4 65 (+65%) 71.3 (+81%) 96.3 (+144%) 59.9 57.5 56.3 (1.1x) 51.1 (1.2x) 30.7 (1.9x)
@@ IV-B Stage II: Online Fine-tuning with HIL-ConRFT
$\mathcal{L}_{\pi}^{online}(\psi)=\beta\mathcal{L}_{\pi}^{BC}+\eta\mathcal{L}_{\pi}^{Q}$ (5)
where BC loss $\mathcal{L}_{\pi}^{BC}=\mathbb{E}_{(s,a)\sim(\mathcal{D}\cup\mathcal{R}),m\sim\mathcal{U}[1,M-1]}[d(f_{\psi}(a+k_{m}z,k_{m}|E(s)),a)]$, $z\sim\mathcal{N}(0,I)$, $d$ stands for the Euclidean distance $d(x,y)=||x-y||_{2}$, and Q loss $\mathcal{L}_{\pi}^{Q}=-\mathbb{E}_{s\sim(\mathcal{D}\cup\mathcal{R}),a\sim\pi_{\psi}}[Q(s,a)]$. Note that this objective closely mirrors Equation 3 from the offline stage, enabling a quick adaption to online fine-tuning.
@@ V-B3 Analysis
Success Rate (%)
Task Kosmos-2(1.6B) PaliGemma(3B)
Pick Banana 60$\rightarrow$100 65$\rightarrow$100
Put Spoon 55$\rightarrow$100 30$\rightarrow$100
Hang Chinese Knot 45$\rightarrow$100 60$\rightarrow$100
Average 53.3$\rightarrow$100 51.7$\rightarrow$100
TABLE IV: Experimental results of ConRFT on different VLA models. We fine-tune RoboVLM [51] with two VLM backbones using our method. Specifically, we fine-tune only the action head while keeping the visual encoders and transformer backbone frozen. All metrics are reported over 20 trials per task.
Practicality of ConRFT across Various VLA Models
ConRFT is highly versatile and can be applied to any VLM-based architecture with an action head. This flexibility stems from its ability to optimize the action generation process independently of the underlying visual encoder, making it adaptable to various VLA frameworks. To further validate its applicability generalization, we test out approach on fine-tuning RoboVLM [51] with two distinct VLM backbones. As shown in Table IV, the results indicate that ConRFT can effectively enhance the performance of various VLAs, improving the success rates across multiple robotic tasks. This ability to fine-tune the action generation while leveraging the pretrained visual components underscores the broad applicability of ConRFT.
@@ VI Limitations
Although our approach demonstrates strong performance and sample efficiency for fine-tuning VLA models in real-world manipulation tasks, several limitations remain.
@@ VI-A Sensitivity to Reward Engineering
In this work, we implement a task-specific binary classifier to calculate the reward for RL. However, the inherent distributional shift between the classifier’s training data and the state-action distributions generated during RL exploration creates a critical vulnerability, as it can lead the learned policy to engage in reward hacking, exploiting unintended behaviors where the classifier provides inaccurate rewards. For instance, the robot might position its end-effector at a specific location that triggers a false positive, causing the policy to converge to an incorrect behavior. Since these reward classifiers typically provide only sparse feedback, the policy may learn slowly, even with the help of human interventions. On the other hand, this reward-driven approach leads to highly specialized policies that are closely tied to the specific conditions of the task, limiting their ability to generalize to new environments. While introducing multi-task dense reward signals could improve sample efficiency and accelerate policy convergence, it would also demand more sophisticated reward engineering for real-world applications.